# An extended symmetric union with multiple tangle regions and its Alexander polynomial

### v1

@@ 1. Introduction
Let $D$ be a planar diagram of a knot $\hat{K}$ in $S^{3}$. We consider the diagram $D^{\ast}$ as the diagram of the knot obtained by reflecting $\hat{K}$ across the plane orthogonal to the plane on which the diagram $D$ lies. Let $T_{i}$ be tangles such that the left-side endpoints are connected within $T_{i}$, and the right-side endpoints are also connected within $T_{i}$. We call this type of tangle an “even type” tangle. We take $n+1$ arcs $\tilde{x}_{0},\dots,\tilde{x}_{n-1},\tilde{x}_{n}$ from the diaglam $D$ as shown on the left of Figure 1. Let $\tilde{x}_{0}^{\ast},\dots,\tilde{x}_{n-1}^{\ast},\tilde{x}_{n}^{\ast}$ be arcs in $D^{\ast}$ corresponding to $\tilde{x}_{0},\dots,\tilde{x}_{n-1},\tilde{x}_{n}$. Placing the diagrams $D$ and $D^{\ast}$ symmetrically, forming a connected sum of $D$ and $D^{\ast}$ by using $\tilde{x}_{0}$ and $\tilde{x}_{0}^{\ast}$, and replacing each pair of arcs $\tilde{x}_{i}$ and $\tilde{x}_{i}^{\ast}$ with the tangle $T_{i}$, we define a knot $K$ as consisting of tangle regions $\{T_{i}\}$ along with the diagrams $D$ and $D^{\ast}$ as shown on the right of Figure 1.
@@
As a corollary, similar to [8, Corollary 1.5], we can easily construct a non-fibered knot $K$ whose knot group surjectively maps to that of a fibered knot $\hat{K}$, and the longitude of $K$ is mapped to a trivial element by applying the construction from Theorem 1.1 as follows. Let $\hat{K}$ be a fibered knot; for example, we could take $\hat{K}$ to be the figure eight knot. Also, let $T_{k}$ be a tangle for which the Alexander polynomial of numerator $N(T_{k})$ is not monic. Then we see that the Alexander polynomial of $K$, constructed using the method of Theorem 1.1 and including the tangle $T_{k}$ in the set of tangles $\{T_{i}\}$, is not monic. Therefore, the constructed knot $K$ is not fibered since a fibered knot has a monic Alexander polynomial.
@@ 2. Proof of Main Theorem
If there is a region that does not touch the crossings, then a trivial link component exists, and consequently $\Delta_{L}(t)=0$. Therefore, we assume that all regions touch some crossings, which implies that each edge in the diagram of $L$ connects two crossing points, and each crossing point has four edges.
@@
To calculate the Alexander matrix of $L$, four labels $x$, x-x, 11, and 1-1 are assigned to the four regions around a crossing point $d_{i}$ and $d_{i}^{\ast}$ as shown in Figure 4. Figure 4 also shows the correspondence between the labels of regions $c_{i}$ and $c_{i}^{\ast}$ around crossing points $d_{i}$ and $d_{i}^{\ast}$, respectively.
@@
We shall prove Theorem1.1(1) by induction on the number $n$ of tangle regions. Let $K_{n}$ be the knot with $n$ tangle regions described in Theorem 1.1. The case $n=1$ has already been proved in [8]. We assume that $K_{n-1}$ satisfies $\Delta_{K_{n-1}}(t)=\Delta_{N(T_{1})}(t)\cdots\Delta_{N(T_{n-1})}(t)\left(\Delta_{\hat{K}}(t)\right)^{2}$, which is equivalent to $\nabla_{K_{n-1}}(z)=\nabla_{N(T_{1})}(z)\cdots\nabla_{N(T_{n-1})}(z)\left(\nabla_{\hat{K}}(z)\right)^{2}$, where $\nabla_{K}(z)$ denotes the Conway polynomial of a knot $K$. We will prove that $\nabla_{K_{n}}(z)=\nabla_{N(T_{n})}(z)\nabla_{K_{n-1}}(z)$. The proof described below is an analog of Conway’s fraction formula proof (see [2, Theorem 7.9.1], [6, Theorem 2.3]).
@@
Let $\gamma$ be a concatenation of two words $\gamma_{-}$ and $\gamma_{+}$, such that $\gamma=\gamma_{-}\cdot\gamma_{+}$. According to the previous construction, a loop represented by $\gamma$ is parallel to the knot $K$ and the linking number between the loop and $K$ is equal to zero. Therefore, $\gamma$ represents the longitude $\lambda_{K}$ of $K$.
We will precisely describe the structure of sub-words in $\gamma$ as follows. On the both sides of point $p$, at the first encountered under-crossing points $p_{1}$ and $p_{1}^{\ast}$, the added words $w_{1}$ and $w_{1}^{\ast}$ included in the diagrams $D$ and $D^{\ast}$ respectively are of the following types: when the under-corssing point $p_{1}$ is positive, $(w_{1},w_{1}^{\ast})=(x_{i_{1}}^{-1}y_{i_{1}},(y_{i_{1}}^{\ast})^{-1}x_{i_{1}}^{\ast})$, or when $p_{1}$ is negative, $(w_{1},w_{1}^{\ast})=(x_{i_{1}}y_{i_{1}}^{-1},y_{i_{1}}^{\ast}(x_{i_{1}}^{\ast})^{-1})$ as shown in Figure 7.
@@
We next consider the knot groups G(K^)G($\hat${K}) and $G(K)$. The knot group G(K^)G($\hat${K}) is generated by the arcs $\hat{x}_{1},\dots,\hat{x}_{m}$, which correspond to the arcs $x_{1},\dots,x_{m}$ in the diagram $D$. The knot group $G(K)$ is also generated by the arcs $x_{1},\dots,x_{m},x_{1}^{\ast},\dots,x_{m}^{\ast}$ in the diagrams $D$ and $D^{\ast}$, $\alpha_{0}$, $\alpha_{1}$, and $s_{k}^{i}$ in the tangles $T_{i}$. We regard these generators as those of the Wirtinger presentation, and let $r_{i}$ be a rerator corresponding to a crossing in $D$, $r_{i}^{\ast}$ to a crossing in $D^{\ast}$, and $r_{k}^{i}$ to a crossing in $T_{i}$. Note that the generators corresponding to the arcs $\tilde{x}_{0}$ and $\tilde{x}_{0}^{\ast}$ in the diagram $D$ and $D^{\ast}$ are devided and transformed into generators corresponds to the arcs $\alpha_{0}$ and $\alpha_{1}$, similarly $\tilde{x}_{i}$ and $\tilde{x}_{i}^{\ast}$ have also been transformed into $s_{k}^{i}$, $k=1,2,3,4$.
@@ 3. Examples
The knot $K_{A}$ has a single tangle region, so this type of construction has already been discussed in [8]. By adding the tangle $T_{3}^{KT}$ at the position of the twice twisted part of $T_{3}^{KT}$ and $(T_{3}^{KT})^{\ast}$ in $K_{A}$, we obtain a knot $K_{B}$ on the left of Figure 11. We can also see by using SnapPy that the Alexander polynomial is trivial $\Delta_{K_{B}}(t)=1$, and the Jones polynomial is $V_{K_{B}}=t^{-6}-3t^{-5}+5t^{-4}-5t^{-3}+t^{-2}+5t^{-1}-10+14t-13t^{2}+9t^{3}-2t^{4}-5t^{5}+10t^{6}-11t^{7}+8t^{8}-4t^{9}+t^{10}$.
@@
Horie, Kitano, Matsumoto, and Suzuki provided a complete list of pairs $K_{1}\geq K_{2}$ for cases where $K_{1}$ has 11 crossings and $K_{2}=3_{1},4_{1},5_{1},5_{2},6_{1},6_{2},6_{3}$ [4]. The three knots 11a157, 11a264, and 11a305 are the only knots on the list for the case $K_{2}=3_{1}$ where the Alexander polynomial is divisible by the square of the Alexander polynomial of the trefoil and their quotients are products of two polynomials. The knot 11a157 can be constructed using our construction for the case of two tangle regions, as shown on the left of Figure 12, which is denoted by $K_{D}$. The knot $K_{D}^{\prime}$ on the right of Figure 12 is formed by taking a single tangle region as the connected sum of the trefoil and the figure eight knot. However, the knot $K_{D}^{\prime}$ is also recognized as 11a157 by SnapPy. Although the Alexander polynomials of the other knots 11a264 and 11a305 are the same as those of 11a157, we have not yet succeeded in constructing these knots as knots built using our construction.
@@
Next, we examine the knots in the list that have an epimorphism to the knot group of the figure eight knot. The knots in the list whose Alexander polynomials are divisible by the square of the Alexander polynomial of the figure eight knot are 11a5 and 11a297. The Alexander polynomials of these knots are as follows.
@@
Determine whether the knots 11a264 and 11a305 can be constructed using the method of Theorem 1.1 in the case that the trefoil is a partial knot. Also, determine this for 11a297 in the case that the figure eight knot is a partial knot.

### v2

@@ 1. Introduction
Let $D$ be a planar diagram of a knot $\hat{K}$ in $S^{3}$. We consider the diagram $D^{\ast}$ as the diagram of the knot obtained by reflecting $\hat{K}$ across the plane orthogonal to the plane on which the diagram $D$ lies. Let $T_{i}$ be tangles such that the left-side endpoints are connected within $T_{i}$, and the right-side endpoints are also connected within $T_{i}$. We call this type of tangle an “even type” tangle. We take $n+1$ arcs $\tilde{x}_{0},\dots,\tilde{x}_{n-1},\tilde{x}_{n}$ from the diagram $D$ as shown on the left of Figure 1. Let $\tilde{x}_{0}^{\ast},\dots,\tilde{x}_{n-1}^{\ast},\tilde{x}_{n}^{\ast}$ be arcs in $D^{\ast}$ corresponding to $\tilde{x}_{0},\dots,\tilde{x}_{n-1},\tilde{x}_{n}$. Placing the diagrams $D$ and $D^{\ast}$ symmetrically, forming a connected sum of $D$ and $D^{\ast}$ by using $\tilde{x}_{0}$ and $\tilde{x}_{0}^{\ast}$, and replacing each pair of arcs $\tilde{x}_{i}$ and $\tilde{x}_{i}^{\ast}$ with the tangle $T_{i}$, we define a knot $K$ as consisting of tangle regions $\{T_{i}\}$ along with the diagrams $D$ and $D^{\ast}$ as shown on the right of Figure 1.
@@
As a corollary, similar to [8, Corollary 1.5], we can easily construct a non-fibered knot $K$ whose knot group surjectively maps to that of a fibered knot $\hat{K}$, and the longitude of $K$ is mapped to a trivial element by applying the construction from Theorem 1.1 as follows. Let $\hat{K}$ be a fibered knot; for example, we could take $\hat{K}$ to be the figure-eight knot. Also, let $T_{k}$ be a tangle for which the Alexander polynomial of numerator $N(T_{k})$ is not monic. Then we see that the Alexander polynomial of $K$, constructed using the method of Theorem 1.1 and including the tangle $T_{k}$ in the set of tangles $\{T_{i}\}$, is not monic. Therefore, the constructed knot $K$ is not fibered since a fibered knot has a monic Alexander polynomial.
@@ 2. Proof of Main Theorem
If there is a region that does not touch any of the crossings, then a trivial link component exists, and consequently $\Delta_{L}(t)=0$. Therefore, we assume that all regions touch some crossings, which implies that each edge in the diagram of $L$ connects two crossing points, and each crossing point has four edges.
@@
To calculate the Alexander matrix of $L$, four labels $x$, x-x, 11, and 1-1 are assigned to the four regions around a crossing point $d_{i}$ and $d_{i}^{\ast}$ as shown in Figure 4. Figure 4 also shows the correspondence between the labels of regions $c_{i}$ and $c_{i}^{\ast}$ around the crossing points $d_{i}$ and $d_{i}^{\ast}$, respectively.
@@
We shall prove Theorem1.1(1) by induction on the number $n$ of tangle regions. Let $K_{n}$ be the knot with $n$ tangle regions described in Theorem 1.1.
In the case of $n=1$, the diagram of $K_{1}$ is regarded as the one shown in Figure 5, where $\tilde{T}_{1}$ is the tangle obtained by rotating $T_{1}$ by $\pi$. Let $T_{0}$ be the tangle formed by connecting two diagrams $D$ and $D^{\ast}$.
Figure 5. the diagram of $K_{1}$ in the case $n=1$
The tangle sum of two tangles $\tilde{T}_{1}$ and $T_{0}$ denotes $\tilde{T}_{1}+T_{0}$, and we observe that $K_{1}=N(\tilde{T}_{1}+T_{0})$ according to the notation. $\nabla_{K}(z)$ denotes the Conway polynomial of $K$, and $D(T)$ denotes the denominator of the tangle $T$ which is constructed by connecting the left-side endpoints of $T$ to each other and by also connecting the right-side endpoints. Applying Conway’s fraction formula (see, [2, Theorem 7.9.1], [6, Theorem 2.3]) to $\nabla_{N(\tilde{T}_{1}+T_{0})}$, we obtain the following formula;
$\nabla_{N(\tilde{T}_{1}+T_{0})}(z)=\nabla_{N(\tilde{T}_{1})}(z)\nabla_{D(T_{0})}(z)+\nabla_{D(\tilde{T}_{1})}(z)\nabla_{N(T_{0})}(z).$
$N(T_{0})$ is equivalent to the $L_{0}^{(1)}$ that appears in Lemma 2.1 for the case $n=1$. Then $\nabla_{N(T_{0})}(z)=0$ because $\Delta_{N(T_{0})}(t)=0$. Furthermore, since $D(T_{0})$ is equivalent to the connected sum of two diagrams $D$ and $D^{\ast}$, we see that $\nabla_{D(T_{0})}(z)=\left(\nabla_{\hat{K}}(z)\right)^{2}$. By the above argument and the fact that $N(\tilde{T}_{1})$ is equivalent to $N(T_{1})$, we can see the following;
$\nabla_{K_{1}}(z)=\nabla_{N(\tilde{T}_{1}+T_{0})}(z)=\nabla_{N(T_{1})}(z)\left(\nabla_{\hat{K}}(z)\right)^{2},$
and we conclude that $\Delta_{K_{1}}(t)=\Delta_{N(T_{1})}(t)\left(\Delta_{\hat{K}}(t)\right)^{2}$. This completes the proof for the case of $n=1$.
We next assume that $K_{n-1}$ satisfies $\Delta_{K_{n-1}}(t)=\Delta_{N(T_{1})}(t)\cdots\Delta_{N(T_{n-1})}(t)\left(\Delta_{\hat{K}}(t)\right)^{2}$, which is equivalent to $\nabla_{K_{n-1}}(z)=\nabla_{N(T_{1})}(z)\cdots\nabla_{N(T_{n-1})}(z)\left(\nabla_{\hat{K}}(z)\right)^{2}$. We will prove that $\nabla_{K_{n}}(z)=\nabla_{N(T_{n})}(z)\nabla_{K_{n-1}}(z)$. The proof described below is an analog of Conway’s fraction formula proof.
@@
Let $\gamma$ be a concatenation of two words $\gamma_{-}$ and $\gamma_{+}$, such that $\gamma=\gamma_{-}\cdot\gamma_{+}$. According to the previous construction, a loop represented by $\gamma$ is parallel to the knot $K$, and the linking number between the loop and $K$ is equal to zero. Therefore, $\gamma$ represents the longitude $\lambda_{K}$ of $K$.
We will precisely describe the structure of sub-words in $\gamma$ as follows. On both sides of point $p$, at the first encountered under-crossing points $p_{1}$ and $p_{1}^{\ast}$, the added words $w_{1}$ and $w_{1}^{\ast}$ included in the diagrams $D$ and $D^{\ast}$ respectively are of the following types: when the under-crossing point $p_{1}$ is positive, $(w_{1},w_{1}^{\ast})=(x_{i_{1}}^{-1}y_{i_{1}},(y_{i_{1}}^{\ast})^{-1}x_{i_{1}}^{\ast})$, or when $p_{1}$ is negative, $(w_{1},w_{1}^{\ast})=(x_{i_{1}}y_{i_{1}}^{-1},y_{i_{1}}^{\ast}(x_{i_{1}}^{\ast})^{-1})$ as shown in Figure 8.
@@
We next consider the knot groups G(K^)G($\hat${K}) and $G(K)$. The knot group G(K^)G($\hat${K}) is generated by the arcs $\hat{x}_{1},\dots,\hat{x}_{m}$, which correspond to the arcs $x_{1},\dots,x_{m}$ in the diagram $D$. The knot group $G(K)$ is also generated by the arcs $x_{1},\dots,x_{m},x_{1}^{\ast},\dots,x_{m}^{\ast}$ in the diagrams $D$ and $D^{\ast}$, $\alpha_{0}$, $\alpha_{1}$, and $s_{k}^{i}$ in the tangles $T_{i}$. We regard these generators as those of the Wirtinger presentation, and let $r_{i}$ be a relator corresponding to a crossing in $D$, $r_{i}^{\ast}$ to a crossing in $D^{\ast}$, and $r_{k}^{i}$ to a crossing in $T_{i}$. Note that the generators corresponding to the arcs $\tilde{x}_{0}$ and $\tilde{x}_{0}^{\ast}$ in the diagram $D$ and $D^{\ast}$ are devided and transformed into generators corresponding to the arcs $\alpha_{0}$ and $\alpha_{1}$, similarly $\tilde{x}_{i}$ and $\tilde{x}_{i}^{\ast}$ have also been transformed into $s_{k}^{i}$, $k=1,2,3,4$.
@@ 3. Examples
The knot $K_{A}$ has a single tangle region, so this type of construction has already been discussed in [8]. By adding the tangle $T_{3}^{KT}$ at the position of the twice twisted part of $T_{3}^{KT}$ and $(T_{3}^{KT})^{\ast}$ in $K_{A}$, we obtain a knot $K_{B}$ on the left of Figure 12. We can also verify with SnapPy that the Alexander polynomial is trivial $\Delta_{K_{B}}(t)=1$, and the Jones polynomial is $V_{K_{B}}=t^{-6}-3t^{-5}+5t^{-4}-5t^{-3}+t^{-2}+5t^{-1}-10+14t-13t^{2}+9t^{3}-2t^{4}-5t^{5}+10t^{6}-11t^{7}+8t^{8}-4t^{9}+t^{10}$.
@@
Horie, Kitano, Matsumoto, and Suzuki provided a complete list of pairs $K_{1}\geq K_{2}$ for cases where $K_{1}$ has 11 crossings and $K_{2}=3_{1},4_{1},5_{1},5_{2},6_{1},6_{2},6_{3}$ [4]. The three knots 11a157, 11a264, and 11a305 are the only knots on the list for the case $K_{2}=3_{1}$ where the Alexander polynomial is divisible by the square of the Alexander polynomial of the trefoil and their quotients are products of two polynomials. The knot 11a157 can be constructed using our construction for the case of two tangle regions, as shown on the left of Figure 13, which is denoted by $K_{D}$. The knot $K_{D}^{\prime}$ on the right of Figure 13 is formed by taking a single tangle region as the connected sum of the trefoil and the figure-eight knot. However, the knot $K_{D}^{\prime}$ is also recognized as 11a157 by SnapPy. Although the Alexander polynomials of the other knots 11a264 and 11a305 are the same as those of 11a157, we have not yet succeeded in constructing these knots as knots built using our construction.
@@
Next, we examine the knots in the list that have an epimorphism to the knot group of the figure-eight knot. The knots in the list whose Alexander polynomials are divisible by the square of the Alexander polynomial of the figure-eight knot are 11a5 and 11a297. The Alexander polynomials of these knots are as follows.
@@
Determine whether the knots 11a264 and 11a305 can be constructed using the method of Theorem 1.1 in the case that the trefoil is a partial knot. Also, determine this for 11a297 in the case that the figure-eight knot is a partial knot.